News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 6 August 1999 7.00am EST

The following is an overview of the events that occurred over the last 24 hours. Measured optical cavity pressure of 8.3 E-06 Torr. Pressure has been reduced from 8.6 E-06 in the last 24 hours.

Continued ACIS calibrations.

Continued development of burn strategy and command loads for IPS #5.

Continued Earth Scans to update Chandra attitude reference accuracy.

Events scheduled for the next 24 hours
  • Continue ACIS and HRC calibrations.
  • Perform IPS #5 burn, retaining the current apogee of 139,141km. Perigee will be raised by this burn to an altitude within the range of 9,000kn to 10,000km which minimizes the effects of lunar eclipse on the Chandra electrical power system.
  • Perform post burn propellant slosh evaluation to assure no adverse effect on PCAD performance.
  • Evaluate Aspect Camera performance sensitivity as a function of HRMA thermal control system operation.






  • Return to Status Reports